# Hydrodynamic noise in one dimension: projected Kubo formula and its vanishing in integrable models

### v1

@@ Appendix B Drude weight from nonlinear fluctuating Boltzmann-Gibbs principle
It is instructive to verify that our formalism is in agreement with the general result (114), by evaluating the Drude weight using the nonlinear fluctuating Boltzmann-Gibbs principle (3.1).
@@
Consider $\langle\!\langle o_{1}(x_{1},t_{1}),o_{2}(x_{2},t_{2})\rangle\!\rangle^{\rm c}$, which depends on $x_{1}-x_{2}$ and $t_{1}-t_{2}$ by space-time translation invariance. We may integrate over $x_{1}$, or integrate over $x_{2}$, and we look at the limit $t_{1}-t_{2}\to\pm\infty$. Clearly, the noise-boise term (70) gives zero under the limit $t_{1}-t_{2}\to\pm\infty$, and by the results of Sec. 4.3 all other correlations vanish except for that from the microcanonical terms in (3.1), that is (4.3).
For the fist line on the right-hand side of (4.3), we obtain (114). Hence, we need to verify that the other lines give zero.

### v2

@@ Appendix D Drude weight from nonlinear fluctuating Boltzmann-Gibbs principle
It is instructive to verify that our formalism is in agreement with the general result (197), by evaluating the Drude weight using the nonlinear fluctuating Boltzmann-Gibbs principle (21).
@@
Consider $\langle\!\langle o_{1}(x_{1},t_{1}),o_{2}(x_{2},t_{2})\rangle\!\rangle^{\rm c}$, which depends on $x_{1}-x_{2}$ and $t_{1}-t_{2}$ by space-time translation invariance. We may integrate over $x_{1}$, or integrate over $x_{2}$, and we look at the limit $t_{1}-t_{2}\to\pm\infty$. Clearly, the noise-boise term (71) gives zero under the limit $t_{1}-t_{2}\to\pm\infty$, and by the results of Sec. 4.2 all other correlations vanish except for that from the microcanonical terms in (21), that is (4.2).
For the first line on the right-hand side of (4.2), we obtain (197). Hence, we need to verify that the other lines give zero.